In order to financially stimulate the nation, the Federal government injected $900 billion dollars into the economy. However, the results were less than spectacular. One reason could have been a failure to understand the marginal propensity to consume. Assume the marginal propensity to consume (MPC) was only 0.4. How much of that $900 billion went to increased consumption? Where did the rest of the money go?

Increased consumption: _______________________

Where did the rest go? _____________________

Using MPC = 0.4, what is the spending multiplier (the actual numerical value please): ______________

What was the overall change in income as a result of the stimulus package after the multiplier completely works its way through the economy? ___ __________
Answer #1

ニリ(1-04) = 1(06-10/6-1, 66 aine Change in overal io Come in the ee economy would be multi ye云米90 oaq hu
Answer #2

Ans:

Increased consumption = mpc*$900 billion = 0.4*900 billion = $360 billion

The rest $540 billion were part of savings which further translated to investments..

spending multipier = 1/1-mpc = 1/0.6 = 1.667

After multiplier completely worked, $360 billion of government investment which flowed through consumption would have translated to 1.667*$900 billion as multiplier effect comes in picture . So, overall change in income is $1500.3 billions.